### v3

# IMAIA: Interactive Maps AI Assistant for Travel Planning and Geo-Spatial Intelligence

Abstract
Map applications are still largely point-and-click, making it difficult to ask map-centric questions or connect what a camera sees to the surrounding geospatial context with view-conditioned. We introduce IMAIA, an interactive Maps AI Assistant that enables natural-language interaction with both vector (street) maps and satellite imagery, and augments camera inputs with geospatial intelligence to help users understand the world. IMAIA comprises two complementary components. Maps Plus treats the map as first-class context by parsing tiled vector/satellite views into a grid-aligned representation that a language model can query to resolve deictic references (e.g., “the flower-shaped building next to the park in the top-right”). Places AI Smart Assistant (PAISA) performs camera-aware place understanding by fusing image–place embeddings with geospatial signals (location, heading, proximity) to ground a scene, surface salient attributes, and generate concise explanations. A lightweight multi-agent design keeps latency low and exposes interpretable intermediate decisions. Across map-centric QA and camera-to-place grounding tasks, IMAIA improves accuracy and responsiveness over strong baselines while remaining practical for user-facing deployments. By unifying language, maps, and geospatial cues, IMAIA moves beyond scripted tools toward conversational mapping that is both spatially grounded and broadly usable.
I Introduction
Modern map applications remain largely point–and–click: users pan and zoom, then issue basic, limited and inflexible queries. This interaction model breaks down for map-centric, view-conditioned queries—e.g., “What’s the flower-shaped building next to the park in the top-right of what I’m viewing?”—and for connecting what a mobile camera sees to the surrounding geospatial context. The desire to explore and navigate unfamiliar environments is fundamental, yet current tools remain inadequate: traditional travel and mapping systems, constrained by static methodologies, struggle with real-world dynamism—fluctuating conditions, imprecise signals, and unexpected disruptions that degrade the experience. Meanwhile, travel planning, navigation, and local discovery are typically engineered as isolated modules, yielding fragmented interactions and brittle handoffs [1, 2]; prior work even documents user behavior under disrupted plans [3], yet practical systems still fail to support coherent, conversational map understanding.
The recent rapid advancements in Large Language Models (LLMs) present a transformative opportunity to transcend these limitations [4, 5, 6]. LLMs possess an unprecedented ability to process and synthesize diverse multimodal inputs, including text, imagery, geospatial data, and contextual cues [7]. This capability is paving the way for a new generation of intelligent systems that are not only cohesive but also inherently adaptive [8]. This work is situated at the confluence of two pivotal trends in geospatial AI. Firstly, LLMs are increasingly adept at interpreting unstructured or ambiguous geospatial information, transforming vague user requests into precise map coordinates or deriving actionable insights from noisy datasets [8]. Secondly, advancements of conversational AI underscores the necessity of multi-turn, context-aware interactions for complex tasks such as navigation and discovery, where user needs and intentions can shift dynamically [9]. Beyond this, Vision–language models (VLMs [10]) can describe images, but without explicit spatial grounding to the current map state (viewport, scale, nearby entities) and geospatial signals (location, heading, proximity), responses are often brittle or slow. Rather than fragmenting planning, navigation, and local discovery into separate and loosely coupled modules, our approach unifies these capabilities within a single cohesive framework.
In this paper, we present the Interactive Maps AI Assistant (IMAIA), depicted in Figure 1. IMAIA is built around two tightly integrated components—Maps Plus and the Places AI Smart Assistantvthat operate in concert under a multi-agent orchestration layer. By coupling Maps Plus with PAISA, the system eliminates cross-module discontinuities and enables seamless reasoning across both map-centric spatial analysis and camera-based place understanding. This design supports fluid transitions between high-level reasoning on maps and fine-grained perception of real-world scenes, resulting in a coherent end-to-end user experience. The overarching goal of IMAIA is to deliver an interactive geospatial platform that bridges online trip planning with on-site exploration and discovery. By leveraging the complementary strengths of large language models (LLMs), vision–language models (VLMs), and a flexible multi-agent architecture, IMAIA provides a mapping experience that is not only more natural and adaptive but also capable of intelligently supporting both digital exploration and real-world navigation.
Maps Plus (map-centric grounding). We parse vector/satellite maps into a grid-aligned representation and align detected entities (e.g., roads, parks, water bodies) to a geospatial index, enabling efficient resolution of deictic, view-conditioned queries (e.g., “the lake in the top-right tile”) that support intelligent trip planning.
PAISA (multimodal understanding). PAISA fuses visual input from the camera with geospatial signals—including location, heading, and proximity—within a coordinated multi-agent architecture. By combining these heterogeneous modalities across orchestration, location intelligence, navigation, and spatial reasoning agents, the system delivers guidance that is both concise and context-aware, grounding real-world assistance in a rich multimodal understanding of the user’s environment.
IMAIA as a unified system. By tightly integrating Maps Plus and PAISA within a coordinated framework, IMAIA establishes an end-to-end pipeline that bridges map-based exploration with environment-grounded place understanding. Empirically, IMAIA achieves substantial improvements over strong baselines. Maps Plus raises place detection accuracy from under 43% to nearly 90%, while PAISA enables human-centered navigation that reduces detours compared to rigid turn-by-turn instructions. For spatial reasoning, our distilled model reaches 84% accuracy—a 3× gain over large multimodal LLMs—and delivers a 7.3 inference speedup (1.7s vs. 12.4s) over agent-based pipelines, ensuring responsiveness in real-world settings.
II Related Work
II-A Geospatial/Maps Intelligence with LLMs
Recent advancements in Large Language Models (LLMs) [11, 12] are driving a paradigm shift in information retrieval, moving from single-query, text-only systems to conversational, multi-modal search. This trend is particularly prominent in the geospatial domain, where models have been empowered with grounded language understanding [13]. A growing body of work has demonstrated the potential of LLMs for enriched map searches, adeptly handling spatio-temporal data and conversational queries [5, 8, 14, 15, 16]. However, a key challenge remains largely unaddressed: how to efficiently feed information from existing geo-indexing systems into an LLM. Little research has focused on this critical interface, which is necessary for the performant resolution of complex, multi-modal geospatial queries.
II-B Spatial Intelligence and Reasoning with VLMs
Recent works have attempted to augment vision–language models (VLMs) with spatial reasoning capabilities. ASMv2 [17] introduces fine-tuned modules for spatial VQA, SpatialVLM [18] synthesizes large-scale spatial question–answer pairs to improve metric distance estimation, and SpatialRGPT [19] incorporates scene graphs for relational reasoning. While these approaches advance spatial intelligence, they often exhibit two key limitations: (i) high latency, due to reliance on computationally intensive synthetic pipelines or graph-based reasoning, and (ii) task misalignment, as training objectives are primarily benchmark-driven rather than optimized for embodied agent tasks.
III Interactive Maps AI Assistant
III-A Maps Plus
Travelers usually begin by asking broad questions such as “What are the top locations to visit in X?”, “What should I see in X?”, or “Plan me a route through X.” Previous studies have shown that a large language model (LLM) can answer such queries either directly from its internal memory or by coordinating with external tools [8, 5], as illustrated in Figure 3. These approaches are referred to as Single Model, Model + Location, and Model + Verbose Location. Once users begin examining an interactive map, however, they often pose richer, map-centric questions that require spatial reasoning about what they are actually viewing rather than a static list of attractions. To meet these needs, we introduce a multimodal system that blends an LLM with image input and geospatial search (User Interface as shown in Figure 2). A user can click on map tiles or satellite imagery and converse naturally about that view, receiving answers grounded in both textual knowledge and visual context. As shown in Figure 3, our proposed method first determines the geographic focus and zoom level of the user’s current view, then scans the surrounding imagery on a regular grid to extract visual features and detect salient geographic entities. Finally, it queries a geospatial index with those entities and synthesizes the results so the LLM can craft an informed, location-aware response. By treating the map itself as conversational context, the system supports fluid trip planning and exploratory tasks that go well beyond what text-only approaches can deliver.
III-A1 Location Awareness
The first step involves providing GPT-4o with contextual information about where an image was taken. This can be presented in a structured format, such as precise latitude and longitude coordinates (e.g., 42.344, 36.236) or as a verbose description of the place (e.g., Seattle, WA, USA). This location-aware capability allows the model to ground its responses in geographic context, ensuring more relevant and accurate interpretations of map-based queries.
III-A2 Grid-Based Spatial Analysis
Next, a simplified map with a grid overlay is provided to multimodal LLM as shown in Figure 4. The model is tasked with identifying grid cells that contain significant map entities, such as roads, parks, or water bodies. This step allows for spatial correlation analysis by associating detected entities with their positions on the map. By segmenting the map in this way, the model can break down complex spatial relationships and make them more accessible for downstream tasks, such as answering questions about specific regions.
III-A3 Entity Search and Query Resolution
When a user asks a question such as ”What is the lake at the top right part of the map?”, GPT-4o determines which part of the map they are referring to—such as the ”top right” region—and retrieves relevant geographic entities using the Azure Maps API. The detected entities (e.g., Bonnet Lake, Abi’s Park) are appended to the user’s query and reintroduced to GPT-4o for context-aware reasoning. The model then processes this enhanced prompt and provides a precise answer. By integrating LLM-based reasoning with geospatial search capabilities, this system enables more intuitive interactions with maps, making it a valuable tool for travelers, researchers, and anyone exploring unfamiliar places.
III-B Places AI Smart Assistant (PAISA)
Maps Plus provides a strong foundation for exploring geospatial search capabilities and reasoning. However, it does not offer a natural way to interact with the real world or visually interpret the environment around the user. In practice, when someone is standing in front of an unfamiliar building or exploring new places in an unfamiliar city, relying solely on search-based tools is often insufficient. This gap motivates the need for a system that can seamlessly combine visual understanding with user geolocation, enabling richer, more intuitive, and context-aware interactions. PAISA addresses this challenge by integrating multimodal signals—such as camera input and spatial context—to deliver a more immersive and informative real-world experience. The user interface of PAISA, illustrated in Figure 5, is powered by a backend multi-agent system coordinated by an orchestrator agent. This system incorporates several specialized function agents, including a location intelligence agent, an interactive navigation agent, and a spatial understanding agent. Each agent is powered by an LLM and equipped with a set of functional tools.
III-B1 Interactive Navigation Agent
The Interactive Navigation Agent (INA) is specifically designed to address the last-100-meter problem, helping users navigate the final stretch of their journey with precision. By leveraging the user’s latitude, longitude, orientation, and destination coordinates, the agent guides users through the final segment, ensuring they reach their destination without confusion. The bearing to the destination is calculated using the formula:
Where are latitudes and are longitudes of the user current location and destination location, and is the difference in longitude. Next, the bearing is adjusted for the user’s orientation to find the relative direction (shown in Figure 6):
Finally, ensure the direction is compass-friendly by adjusting for values outside the 0 to 360-degree range. Additionally, INA includes a trigger feature that allows users to view the street view of their destination, offering a visual preview of the surroundings. This functionality enhances the user experience, providing clear and interactive navigation through the most challenging part of the journey, with real-time feedback and immersive, location-based guidance.
III-B2 Location Intelligence Agent
The location intelligence agent enables users to explore unfamiliar places by identifying venues and retrieving relevant information to satisfy their curiosity. As illustrated in Figure 7, this agent first determines what the place is and then leverages metadata and user reviews to enrich the understanding of that location. To ground a user-captured image to the correct venue, we encode the image using a Contrastive Language-Image Pre-training (CLIP) [20] visual encoder, while each candidate place is represented with a CLIP text encoder applied to a structured descriptor that concatenates the place name, category, and latitude/longitude. From these representations, we construct a feature vector consisting of: (i) the cosine similarity between image and place embeddings, (ii) the distance between the user and the place, and (iii) a heading-consistency term defined as the absolute angular difference between the user’s device orientation and the bearing from the user to the place. To further enhance grounding accuracy, we augment this vector with local popularity indicators derived from Azure Maps search activity, providing data-quality priors. The resulting features are fed into an XGBoost ranking model, which assigns relevance scores and reorders the initial retrieval set. The top-ranked candidates are subsequently passed to a downstream LLM agent, supplying a compact, higher-recall context that improves the quality of the final answer.
III-B3 Spatial Understanding Agent
Existing vision–language models fine-tuned for spatial reasoning (e.g., ASMv2 [17], SpatialVLM [18], SpatialRGPT [19]) often suffer from high latency and misalignment with the agent’s task requirements. To overcome these limitations, we distilled a Florence-2 model [21] from GPT-4o [22] with instruction fine-tuning, resulting in a more efficient and task-adapted spatial reasoning module. To further enhance the assistant’s spatial understanding, we introduce a spatial scene-graph understanding agent powered by a fine-tuned lightweight multimodal language model. This agent processes a single photo and identifies the most visually salient objects in the scene—those that are most recognizable or attention-catching (e.g., store signs, flags, building façades)—along with their spatial relationships, expressed either as a scene graph or in natural language. The extracted spatial information can be utilized in two complementary ways: (1) When the assistant retrieves cached street-level images of a destination, the agent generates spatially grounded descriptions that highlight the relative positions of salient objects, helping users more reliably recognize the destination in the real world. (2) When users upload a photo of their current surroundings, the agent analyzes the scene and produces a structured spatial representation, which then serves as contextual input for downstream agents. By grounding both destination imagery and user-provided photos in spatial relationships, the system offers guidance that is more interpretable, robust, and actionable.
The distillation pipeline comprises three stages: (i) entity mining, (ii) object and spatial-relation extraction, and (iii) supervised fine-tuning. In stage (i), we prompt GPT-4o-mini [23] to simulate an urban wayfinding scenario and, given a random street-view image, propose candidate “key” items. Repeating this five times per image over 40,000 street-view images, we retain the top-10 most frequently mentioned items per image as the “key” entities. In stage (ii), we localize each entity’s 2D position using YOLO-World [24] and estimate its 3D depth with Depth Anything V2 [25]. We then assemble a structured record of each entity’s 2D coordinates and distance to the camera, and pair the image with bounding-box overlays and a metric depth map to prompt the GPT-4o for pairwise spatial relationships. Prompts follow the Set-of-Mark Prompting paradigm to enhance visual grounding and reasoning. In stage (iii), we build a supervised fine-tuning dataset by pairing each annotated image with over 15 variants of spatial queries that reflect realistic urban wayfinding needs, such as relative positioning (“What object is directly to the left of the store sign?”) and orientation (“Which structure faces the street?”). Each query is matched with answers derived from the outputs of stage (ii), yielding diverse training samples that encourage the model to generalize beyond simple object detection. We fine-tune a Florence-2-large with full-parameter training, casting the task into its native Dense Region Captioning format so the model learns to generate spatially grounded, linguistically coherent descriptions conditioned on visual regions and relational context. This final stage enables the distilled model to inherit GPT-4o’s reasoning ability while remaining efficient and task-adapted, ultimately producing a lightweight yet robust spatial reasoning module capable of supporting real-time navigation.
III-B4 Handle complex query with multi-agent reasoning
Modern map applications often fall short when handling complex, user-centric queries. Consider a scenario where a user urgently seeks the nearest boba tea shop. Executing this seemingly simple request typically involves multiple steps: initiating a search for boba tea within a specified area, manually reviewing and ranking results based on proximity, and finally selecting a navigation option to begin the route. This fragmented interaction model demands several discrete actions from the user, rather than supporting a seamless, single-command experience—highlighting a common usability limitation in current systems. To address this challenge, we introduce a multi-agent framework capable of interpreting and executing such queries end-to-end. As illustrated in Figure 8, our proposed system, PAISA, directly outputs navigation instructions to the closest boba tea shop based on the user’s location. Naturally, this raises the question of whether such output is a hallucination from a large language model (LLM). To validate the reasoning, we examine PAISA’s internal decision process: the system performs entity search within the target area, computes distances between the user and candidate locations, and ranks them accordingly. In this case, “Boba Express” emerges as the top result due to its proximity—just 1.6 miles from the user—demonstrating the system’s grounded and interpretable reasoning.
The backend of PAISA is organized as a multi-agent system (Figure 9). The orchestrator agent first analyzes the query and decomposes it into simpler sub-queries. These are passed to the location intelligence agent, which retrieves candidate entities and their attributes. The orchestrator then forwards the enriched information to the interactive navigation agent, which generates the optimal route. Finally, the navigation plan is returned to the orchestrator and delivered to the user.
IV Results
IV-A Place detection accuracy with Maps Plus
In this section, we evaluate and compare the quality of geospatial entity search across several methods. Specifically, we benchmark maps plus approach against three widely used baselines (Figure 3) in Table I: (1) Single Model: The model receives only the user query and the current map view image, relying solely on its pretrained knowledge. (2) Model + Location: The input includes the query, the map view image, and the geographic coordinates (latitude and longitude) of the map view. (3) Model with Verbose Location: Similar to the previous setup, but the location input is replaced with verbose descriptors such as city names and landmarks. In this study, we construct a dataset by selecting 10 cities across the United States and sampling points of interest (POIs) within a 20-kilometer radius of each city center. For each POI, we employ GPT-4o to generate synthetic queries based on information such as the POI’s attributes, geographic coordinates, and related contextual data. This process yields a total of 4,300 queries, for example, What is the lake at the top left part of the map.
| Method | Accuracy |
| Single Model | 39.30% |
| Model + Location | 41.46% |
| Model + Verbose Location | 42.74% |
| Maps Plus | 89.83% |
Table I presents the accuracy achieved by various methods. The results indicate that our proposed approach attains significantly higher accuracy compared to the baselines, without requiring any fine-tuning of the LLM. The superior performance of the proposed method can be attributed to its efficient integration of grounding data, specifically entities retrieved from geospatial services, into the LLM’s processing. For instance, consider the query ”What is the coffee shop below the cinema?”. The proposed method not only furnishes a pertinent set of local entities but also deconstructs the query into a series of sub-problems. This multi-step decomposition facilitates improved LLM reasoning and consequently, enhanced performance.
| Ranker | Precision@Top-1 | Recall@Top-1 | Precision@Top-3 | Recall@Top-3 |
| XGBoost Ranker | 80.4% | 72.5% | 36.2% | 92.8% |
| Distance-sorting | 76.1% | 69.2% | 30.4% | 77.5% |
| Similarity-sorting | 65.2% | 58.3% | 25.4% | 68.1% |
IV-B Comparison between human-centered path and turn-by-turn walking directions path
We compare human-centered guidance with conventional turn-by-turn directions routes (Figure 10). Turn-by-turn paths, generated from map topology, are constrained by predefined map data and may introduce unnecessary detours. In practice, pedestrians often prefer more direct or intuitive routes, such as open areas or informal shortcuts, reflecting a gap between algorithmic routing and human spatial reasoning. Human-centered guidance addresses this by pointing users toward the destination, allowing real-time adaptation based on their perception of walkable paths. As illustrated in Figure 11, our system integrates camera, geolocation, and orientation data to compute directional bearings and overlay AR cues. This method reduces detours, increases flexibility, and provides a more natural navigation experience in dynamic or unstructured environments where rigid step-by-step instructions may fail.
IV-C Evaluation on the embedding-based entity search
The XGBoost ranker was trained on a dataset of 500 image queries, each paired with manually annotated ground‑truth venues, and evaluated on a held‑out set of 50 queries. For benchmarking, we considered two baselines: (i) a distance‑based method that orders candidate places solely by geodesic proximity to the user, and (ii) a similarity‑based method that ranks candidates exclusively according to the cosine similarity between image and place embeddings. Table II summarizes the comparative performance of the proposed XGBoost ranker and two baseline methods, evaluated using Precision and Recall at Top‑1 and Top‑3. The Top-k Precision and Recall are defined as:
Top‑1 metrics () measure whether the highest‑ranked candidate is correct, providing a strict indicator of ranking accuracy at the very top. In contrast, Top‑3 metrics () evaluate the proportion and coverage of relevant items within the first three positions, reflecting the system’s ability to surface multiple correct candidates early in the ranking. As shown in Table II, the XGBoost ranker consistently outperforms both distance‑based and similarity‑based baselines across all metrics, with the largest gains observed in Top‑3 recall, indicating improved breadth of relevant retrieval without sacrificing top‑rank precision.
IV-D Evaluation on Spatial Understanding
We evaluate the proposed spatial reasoning module on a test set of 400 street-view images, comparing against both general-purpose multimodal LLMs and specialized spatial-scene graph models. Accuracy is measured with an LLM-as-judge protocol using the OpenAI o1 model, while efficiency and recall are assessed with task-specific metrics. Comparison with multimodal LLMs. As shown in Figure 12, our distilled model achieves an accuracy of 84%, indicating that the majority of generated spatial descriptions are judged as correct by the o1 evaluator. In contrast, Florence-VL 8B [26], a general multimodal LLM built upon Florence-2 with nearly ten times more parameters, attains only 27% accuracy under the same setting. This result highlights the effectiveness of task-aligned distillation for spatial reasoning compared to parameter scaling. Comparison with scene-graph models. Against specialized spatial reasoning systems such as ASM v2 [17], which can generate structured scene graphs but lack natural language interaction capabilities, our model demonstrates higher recall of salient items. On average, our model identifies approximately 7 objects per scene, compared to 4 objects extracted by ASM v2. This improvement suggests that combining structured spatial grounding with natural language reasoning enables richer scene interpretation. Comparison with agent-based solutions. We further benchmark against an agentic pipeline that replicates Stage (i) and (ii) with explicit calls to external models. On a single NVIDIA V100 32GB GPU, the agent-based approach requires 12.4s per image, while our end-to-end distilled model reduces inference time to 1.7s per query, achieving a 7.3 speedup. This efficiency gain is critical for real-time deployment in navigation scenarios. Overall, these results demonstrate that our distilled Florence-2 model achieves strong accuracy, improved recall, and significant efficiency gains over both large-scale multimodal LLMs and agent-based baselines, making it well-suited for spatial reasoning in real-world.
V Conclusion
In this work, we introduced IMAIA, an interactive Maps AI Assistant that unifies language, maps, and geospatial signals to enable natural, spatially grounded interaction. Through its two complementary components—Maps Plus for map-centric question answering and PAISA for camera-aware place understanding—IMAIA significantly improves the accuracy, responsiveness, and usability of geospatial reasoning compared to strong baselines. Our experiments show that Maps Plus boosts place detection accuracy from under 43% with conventional approaches to nearly 90% by treating maps as first-class context and incorporating structured grounding signals. For navigation, PAISA demonstrates that human-centered, direction-based guidance aligns more closely with natural pedestrian behavior than rigid turn-by-turn instructions, offering both flexibility and reduced detours. Beyond place recognition and navigation, our distilled spatial reasoning model achieves 84% accuracy and a 7.3 inference speedup over agent-based pipelines, highlighting the effectiveness of lightweight task-aligned distillation for real-time deployment. Together, these results demonstrate that IMAIA represents a step toward conversational mapping that is both accurate and practical. By integrating visual, spatial, and linguistic cues into a unified framework, it enables richer scene understanding, more intuitive navigation, and scalable performance for user-facing applications.